OpenStudy (anonymous):

Let u = <-3, 4>, v = <8, 2>. Find u + v.

OpenStudy (dan815):

Vector addition, you add the x and y components separetly

OpenStudy (dan815):

u+v=<-3,4>+<8,2>=<-3+8,4+2>=?

OpenStudy (anonymous):

tbh idk how to do this lmao

OpenStudy (dan815):

just addition

OpenStudy (dan815):

<-3+8,4+2>=<____,____>?

OpenStudy (anonymous):

5 , 6

OpenStudy (dan815):